OpenStudy (anonymous):

Does anyone know how to do this ... D=KA[T2-T1/L] and solve for T1

terenzreignz (terenzreignz):

\[\Large D=KA\left[\frac{T_2-T_1}{L}\right]\]

terenzreignz (terenzreignz):

You need to bring everything without \(\large T_1\) to the other side. That means THESE things: \[\Large D=\color{red}{KA}\left[\frac{\color{red}{T_2-}T_1}{\color{red}L}\right]\]

terenzreignz (terenzreignz):

Any idea how?

terenzreignz (terenzreignz):

@Msbballshayla stay with me...

OpenStudy (anonymous):

OH .. no I don't know how to .. I'm still lost at the T2 - T1 over L part

terenzreignz (terenzreignz):

Chill... we're going to use operations on both sides... just remember multiply and divide cancel out add and subtract cancel out

terenzreignz (terenzreignz):

So... the KA bit is multiplied, so how do we remove it?

OpenStudy (anonymous):

Divide it on both sides

terenzreignz (terenzreignz):

That's right. \[\Large \frac{D}{\color{red}{KA}}=\frac{KA\left[\frac{T_2-T_1}{L}\right]}{\color{red}{KA}}\]

terenzreignz (terenzreignz):

These cancel out: \[\Large \frac{D}{\color{}{KA}}=\frac{\cancel{KA}\left[\frac{T_2-T_1}{L}\right]}{\cancel{\color{red}{KA}}}\]

terenzreignz (terenzreignz):

Leaving... \[\Large \frac{D}{KA}= \frac{T_2-T_1}{L}\]

terenzreignz (terenzreignz):

Next step? :)

terenzreignz (terenzreignz):

Let's get rid of the L. How to do that?

OpenStudy (anonymous):

Times L on both side of the equal sign

terenzreignz (terenzreignz):

That's right... I don't see what's troubling you :) \[\Large \color{red}L\times\frac{D}{KA}= \color{red}L\times\frac{T_2-T_1}{L}\]

terenzreignz (terenzreignz):

\[\Large \frac{LD}{KA}= \cancel{L}\times \frac{T_2-T_1}{\cancel{L}}\]

OpenStudy (anonymous):

It was how to get the T2 and T1 from being a fraction

terenzreignz (terenzreignz):

That's already done... L cancels out, see? ^

OpenStudy (anonymous):

Yes I see now.

terenzreignz (terenzreignz):

You get \[\Large \frac{LD}{KA}= T_2-T_1\] Can you do it now? :)

OpenStudy (anonymous):

Do you subtract T2 from both sides now?

terenzreignz (terenzreignz):

Yes, sure, that's possible :) \[\Large \frac{LD}{KA}\color{red}{-T_2}= T_2-T_1 \color{red}{-T_2}\]

terenzreignz (terenzreignz):

Simplified into: \[\Large \frac{LD}{KA}-T_2= -T_1\]Finally?

OpenStudy (anonymous):

I get it now .. but there is still problem similar to this one that still gets me.

terenzreignz (terenzreignz):

Sure. Hit me.

terenzreignz (terenzreignz):

Might want to post a new one though, to avoid lags.

OpenStudy (anonymous):

Okay .. I'll see what problems I have left on my homework to get helped on.